A first order partical differential equations p+q=z-xy is




1
Expert's answer
2021-05-28T02:26:22-0400

This is a Lagrange's equation whose general formula is pP+qQ=R

This can be solved by the formula "(dx\/P)=(dy\/Q)=(dz\/R)"

"Here, P=1 ,Q= 1 ,R=z-xy"

"Now, {dx\/1}={dy\/1}={dz\/z-xy}"

From the 1st two ratio,

dx= dy

=> dx-dy=0

Integrating above equation we get,

"x-y=C \u2015\u2015\u2015(1)"

Now from 1st and 3rd ration,

dx=dz/z-xy

=> "x=ln(z-xy)+C" ――――(2)

Equation (1)&(2) together yeilds the solution.

"x-x+y= ln(z-xy)"

"y= ln(z-xy)"